# Converging to Stability in Two-Sided Bandits: The Case of Unknown Preferences on Both Sides of a Matching Market

### v2

@@ 6.1. Sketch of the Proof of Convergence to Stability Using PCA-UCB
The proof of Theorem 6.1 follows directly from Theorems 6.1 and 6.1. Since the probability of the arms’ confidence intervals overlapping (and thus producing non-deterministic feedback) decays to zero, and because the players’ win estimates converge to their true values with high probability, the overall probability that any “failure” occurs (i.e., either an arm has overlapping intervals or a consequential triplet’s win estimate is inaccurate) tends to zero as time increases. Consequently, the decision rules used by PCA-UCB reduce to those of CA-UCB in the limit.
In summary, our proof strategy first ensures that each arm learns its preference over players via disjoint confidence intervals (Theorem 6.1). This in turn guarantees that the conflict feedback becomes deterministic, allowing the players’ win estimates to converge (Theorem 6.1). Finally, these results imply that the algorithm’s action selection mirrors that of CA-UCB, yielding asymptotic equivalence (Theorem 6.1).
Note that convergence in practice occurs much faster than would be implied by the worst-case analysis. While we derive a bound based around the event $\alpha(T)$, the design of the algorithm ensures that once the confidence intervals for an associated triplet become disjoint, the players immediately receive deterministic feedback—well before the occurrence of $\alpha(T)$. The sooner the players obtain deterministic conflict feedback, the faster the algorithm converges to stability. Moreover, players do not have to wait for $\alpha(T)$ or for the conflict probabilities to converge before acquiring reward estimates for the arms. Consequently, even though the algorithm ultimately replicates CA-UCB, the players begin with some prior information regarding the arms’ rewards, implying that convergence in practice could be significantly tighter, which is consistent with our experimental findings below.
7. Using Thompson Sampling in PCA-SCA (PCA-TS)
Thompson Sampling (Thompson, 1933) (TS) is another approach to solving the MAB problem which has seen a recent resurgence in the literature (Agrawal and Goyal, 2012; Chapelle and Li, 2011; Kong et al., 2022). In the PCA-SCA algorithm, TS can be used to keep track of both the beliefs about rewards by both players and arms. Henceforth, we will refer to this approach as the PCA-TS algorithm. The players and arms keep track of the same information as in UCB, i.e. the total reward obtained for an arm/player, the total number of pulls of an arm, and the total number of conflict wins against each player. The distinction is in how this information is used to update beliefs.
First, players use the same estimates for conflict wins as shown in Equation 4, and subsequently use Equation 1 for an optimistic estimate of winning a conflict. As for the rewards, we assume the variance ($\sigma^{2}=1$) in rewards of the arms and the players are known. Then, we update the mean and precision ($\tau=\frac{1}{\sigma^{2}}$) as:
(5) $\mu_{\text{new}},\tau_{\text{new}}=\frac{\tau_{0}\mu_{0}+\tau\sum_{i=1}^{n}x_{i}}{\tau_{0}+n\tau},\tau_{0}+n\tau$
@@ 7. Using Thompson Sampling in PCA-SCA (PCA-TS)
where $\{\mu_{\text{new}},\tau_{\text{new}}\}$ are the new mean and precision, $\{\mu_{0},\tau_{0}\}$ are the old mean and precision, $\tau=1$ is the known true precision, and $\sum x_{i}$ is the sum of rewards for the agent in question. We can use appropriate lower and upper bounds on the confidence intervals, $\mathcal{R}_{i}^{L}(t),\mathcal{R}_{i}^{U}(t)$.111 In our experiments below, we use $\mathcal{R}_{i}^{L}(t)=\mu_{0}-\frac{1}{\tau_{0}},\mathcal{R}_{i}^{U}(t)=\mu_{0}+\frac{1}{\tau_{0}}$ where $\frac{1}{\tau_{0}}$ represents the standard deviation in belief. These are the lower and upper confidence bounds respectively (Lines 7, 8 respectively in ResolvePullRequests of Algorithm 2). When an agent needs an estimate for a reward for a particular player or an arm, it samples from $\mathcal{N}(\mu_{0},\frac{1}{\tau_{0}})$ where $\mu_{0}$ and $\tau_{0}$ is the corresponding mean and precision tracked by the agent for the specific player/arm. We provide simulation results of using PCA-TS for different market sizes and varied preference heterogeneity in Section 8.
8. Simulation Results
We run simulation experiments where (1) preferences are uniformly random on both sides of the market with varying market sizes $N=K\in\{5,10,15,20\}$; (2) player preference heterogeneity is varied with market size $N=K=10$. The maximum reward an agent can get is $(K+1)$, when matched with its most preferred arm, and the minimum is 1, when matched with the least preferred. The reward gaps (of the means) between consecutively ranked agents are kept constant at $\Delta=1$. To control the level of heterogeneity in player preferences, we use a method due to (Ashlagi et al., 2017), as also used by (Liu et al., 2021). The level of heterogeneity is controlled by a parameter $\beta\in[0,10,100,1000]$. Heterogeneity in preferences decreases with $\beta$, and as $\beta\to\infty$ all players have the same preferences with probability 1. To translate $\beta$ to preference orderings we sample the mean reward $\mu_{i}^{k}$ of arm $a_{k}$ for player $p_{i}$ using the following process:
(7) $\begin{split}x_{k}&\stackrel{{\scriptstyle i.i.d}}{{\sim}}\text{Uniform([0,1])}\\ \epsilon_{k,i}&\stackrel{{\scriptstyle i.i.d}}{{\sim}}\text{Logistic(0, 1)}\\ \overline{\mu}_{i}^{k}&=\beta\cdot x_{k}+\epsilon_{k,i}\\ \mu^{k}_{i}&=\#\{j:\overline{\mu}^{j}_{i}\leq\overline{\mu}_{i}^{k}\}\end{split}$
@@ 8. Simulation Results
These random utilities $\mu_{i}^{k}$ are mapped so that the gap in rewards between consecutively ranked arms is kept constant at $\Delta=1$. The hyper-parameter $\lambda$ is set to $0.9$ and $\kappa$ is set to 10 (in the context of PCA-SCA). [^2] There are alternative methods for controlling heterogeneity in player preferences, for example notions like uncoordinated/coordinated markets (Ackermann et al., 2008). The method we use corresponds to uncoordinated markets when $\beta$ is low. However, coordinated markets are different, with correlation defined on the edges of the matching graph rather than the vertices. We find that the convergence time in these edge-correlated cases is actually comparable to when preferences are uniformly random, compared with node-correlation, where convergence time increases (see Appendix).
@@
8.1. OCA-UCB in the APKP model
Recall that, in this scenario, the players must learn their own position in the arms’ preference order and their own preferences. We compare the performance of this model with the APCK model using CA-UCB. Both models were run with their respective algorithms, with 6000 time steps for random preferences and 3000 time steps for varied preference heterogeneity. The results are shown in Figure 1.
Figure 1. The left sub-figures show the APCK models using CA-UCB and the right sub-figures show the APKP model using OCA-UCB. Results are averaged over 100 runs. The first row varies market size with uniformly random preferences, the second row varies player preference heterogeneity while keeping market size constant at $N=K=10$. OCA-UCB converges to stability even with uncertainty on both sides, though it takes longer (expected because of increased complexity)
The figures in the first row of Figure 1 show results from the experiments where agents’ preferences are uniformly random. We can see that OCA-UCB converges to stability in the APKP model, with increasing market sizes corresponding to slower convergence. This trend is essentially the same as the convergence of CA-UCB in the APCK model, albeit OCA-UCB convergence in the APKP model is slightly slower. Thus, the OCA-UCB algorithm converges to stability in the APKP model almost as fast as the CA-UCB algorithm does in a model where all arm preferences are common knowledge.
The second row in Figure 1 details experimental results for varying player preference heterogeneity. As can be seen from the APCK Model running CA-UCB, decrease in player preference heterogeneity in the APCK model does not have any effect on the convergence of CA-UCB. This trend carries over to OCA-UCB in the APKP model, with convergence slightly slower. This shows that in the setting with varying levels of player preference heterogeneity, OCA-UCB does not show any noticeable difference in the rate at which the markets converge to stability.
In combination with the theoretical results above, our results show that learning to converge to stability is not noticeably harder when arm preferences are not common knowledge, as long as they are known to the arms.
8.2. PCA-UCB and PCA-TS in the APU Model
@@
8.2.1. Using PCA-UCB
The results of the experiments using PCA-UCB in the APU model are summarized in Figure 2. We run PCA-UCB for 20,000 time steps for markets with random preferences on both sides, and for 10,000 time steps for when player preference heterogeneity is varied. The former results are in the first row of the figure, and the results of the latter are detailed in the second row of Figure 2.
@@ 8.2.1. Using PCA-UCB
Figure 2. Results of experiments run in the APU model using UCB to keep track of beliefs: first row has preferences uniformly random, second row has varied player preference heterogeneity with $N=K=10$. Key Takeaways: the markets converge to stability in expectation. If you compare this figure against Figure 3, Thompson Sampling converges to stability slightly faster than UCB and smoother, showing no dependence on player preference heterogeneity.
Firstly, notice that the APU model using PCA-UCB takes longer to stabilize compared to the APKP model using OCA-UCB, as seen in Figures 1 and 2. This is due to the fact that both players and arms need to learn their beliefs from scratch. Regardless, when market preferences are random, the trends in both figures are similar, with market stabilization probabilities approaching 1 and player regret tending to 0. This shows that PCA-UCB can find stable matches in a decentralized fashion with unknown preferences on both sides of the market when preferences are random.
The second row of sub-figures in Figure 2 shows what happens when the players’ preference heterogeneity is varied. We can see that there is no trend associated with $\beta$ concerning the algorithms’ convergence to stability. This is analogous to CA-UCB where a decrease in player preference heterogeneity does not have any effect on the convergence of the algorithm.
8.2.2. Using PCA-TS
Results for when PCA-TS was used in the APU model are shown in Figure 3. Like Figure 2 (the UCB results), the first row in Figure 3 is when the agent preferences are sampled uniformly at random, and the second row is where the players’ preference heterogeneity is varied.
Figure 3. Results of experiments run in the APU model using Thompson Sampling to keep track of beliefs. The first row shows market stability and average player regret when preferences are uniformly random on both sides of the market whereas the second row is for when player preference heterogeneity is varied. All the markets stabilize and player regret tends to 0. Performance was better than UCB in terms of convergence rate and smoothness.
@@ 8.2.2. Using PCA-TS
Results from PCA-UCB (Figure 2) and PCA-TS (Figure 3) show similarities in market stability and player regret for all market sizes with uniformly random preferences, with larger markets stabilizing slower. We can also see that PCA-TS converges faster and more smoothly than PCA-UCB. We will discuss this particular property in further detail in the next subsection. The results demonstrate that Algorithm 2 can find a stable match in the APU model using both UCB and Thompson Sampling.
8.2.3. Convergence Comparison: UCB vs. Thompson Sampling
@@
When looking at Figures 2 and 3, we can see that the Thompson Sampling approach of keeping track of beliefs appears better than UCB in terms of both speed and smoothness of convergence. To quantify this behavior, we turn to the convergence proxy we defined earlier, this time comparing the stability graphs of PCA-UCB and PCA-TS. We set $\mathcal{X}=1000,\theta=90$ and generate these lines, the results of which are summarized in Figure 4. Each color on the graph corresponds to the market size with the solid line representing PCA-UCB and the dotted line representing PCA-TS. We can see that the lines associated with Thompson Sampling reach 1.0 (i.e. 100% of the time-steps in question have stability $>90\%$) quicker and tend to stay that way.
@@
Figure 4. Convergence Comparison of UCB and Thompson Sampling in the APU model with varying market sizes and uniformly random preferences on both sides of the market. The Thompson Sampling approach (dotted) converges slightly faster and stays that way compared to the UCB (solid) which tends to dip time and again before ultimately reaching convergence.
In comparison, the convergence proxy for UCB reaches 1.0 slower for each of the market sizes, and often the proxy value dips before going back up again. By nature of the way beliefs are kept track of, whenever a player chooses to explore, it causes a disturbance in the believed preference ordering of the player. When the internal belief state for a player changes, the player’s set of arms in consideration changes, leading to a different matching. This behavior causes the match to be unstable at some time steps. This is more frequent in the early stages of the algorithm and becomes less frequent as confidence increases about player preferences. PCA-TS exhibits this behavior only for $N=20$ but overall seems to reach 1.0 in a much smoother manner compared to PCA-UCB. The combination of these two reasons suggests that Thompson Sampling might be a better approach for keeping track of agent beliefs in the APU model.
9. Conclusion and Future Work
In this paper, we looked at the problem of two-sided matching with uncertain preferences. While recent work has developed algorithms that converge to a stable match when one side of the market must learn their preferences, we developed a new algorithm that converges to a stable match in two more challenging, settings: the APKP model where the arms have access to their preferences but the agents do not, and the APU model where none of the agents involved know their preferences a priori.
In the APKP model, we prove that the players can reliably learn their positioning in the arms’ preference orderings. This allows the players to construct a set similar to the plausible set from the CA-UCB algorithm and facilitates the convergence of the algorithm to a stable match. We also empirically studied the resulting algorithm, running simulation-based experiments and showing that the matching converges to stability, only slightly slower than the CA-UCB approach in a model where arm preferences are common knowledge. This slowness is expected, because the players now must learn more information before the algorithm can converge.
In the more complex situation represented by the APU model, we propose an algorithm where the players make use of the highest expected reward using an optimism function to pick the arms to attempt at each time step. The players do this by keeping track of estimates of the probability of winning a conflict against another player for each arm to pass to the optimism function in addition to the reward estimates. Then, the players attempt arms that maximize the product of the two. Furthermore, since the arms themselves do not know their preferences in this model, the arms keep track of reward estimates along with the confidence intervals over the players and pick the player with the highest estimated reward from the incoming requests, or if the intervals overlap then the arms pick overlapping players uniformly at random. We theoretically show that the algorithm converges to a stable match when using UCB as the reward estimate. Then we empirically show that the algorithm converges to a stable match, using both UCB and Thomson Sampling as belief estimators where we saw both these approaches converged to stability, with Thompson Sampling approaching stability quicker and more smoothly, mirroring the success of Thompson Sampling in simpler MAB settings.
Prior work has proved convergence in models where arm preferences are common knowledge. In this paper, we have provided new algorithms that provably convergence to stability in both the APKP model and the APU model – the most general setting of the two-sided matching problem.
Appendix A Full Proof of Convergence of PCA-UCB

### v4

@@ 4.1 Using UCB in PCA-SCA (PCA-UCB)
Together, these theorems imply that PCA-UCB not only learns preferences and refines win-estimates efficiently, but ultimately replicates the behavior of the idealized CA-UCB algorithm in the long run.
4.2 Using Thompson Sampling in PCA-SCA (PCA-TS)
Thompson Sampling Thompson (1933) (TS) is an alternative MAB method that has resurged in recent literature Agrawal and Goyal (2012); Chapelle and Li (2011); Kong et al. (2022). In the PCA-SCA algorithm, TS tracks reward beliefs for both players and arms; we call this variant PCA-TS. Like UCB, both players and arms record total rewards, pulls, and conflict wins per player. However, TS updates beliefs differently: players compute conflict win estimates as in Equation 3 and then apply Equation 1 for an optimistic estimate. As for the rewards, we assume the variance ($\sigma^{2}=1$) in rewards of the arms and the players are known. Then, we update the mean and precision ($\tau=\frac{1}{\sigma^{2}}$) as:
$\mu_{\text{new}},\tau_{\text{new}}=\frac{\tau_{0}\mu_{0}+\tau\sum_{i=1}^{n}x_{i}}{\tau_{0}+n\tau},\tau_{0}+n\tau$ (4)
@@ 4.2 Using Thompson Sampling in PCA-SCA (PCA-TS)
where $\{\mu_{\text{new}},\tau_{\text{new}}\}$ are the updated mean and precision, $\{\mu_{0},\tau_{0}\}$ the previous values, $\tau=1$ the known precision, and $\sum x_{i}$ the total reward. We bound the confidence intervals by $\mathcal{R}_{i}^{L}(t)$ and $\mathcal{R}_{i}^{U}(t)$ 111 In our experiments below, we use $\mathcal{R}_{i}^{L}(t)=\mu_{0}-\frac{1}{\tau_{0}},\mathcal{R}_{i}^{U}(t)=\mu_{0}+\frac{1}{\tau_{0}}$ where $\frac{1}{\tau_{0}}$ represents the standard deviation in belief.(see Lines 16 and 17 in ResolvePullRequests of Algorithm 2). When an agent estimates a reward for a player or arm, it samples from $\mathcal{N}(\mu_{0},1/\tau_{0})$ using its tracked $\mu_{0}$ and $\tau_{0}$. Simulation results for PCA-TS under varying market sizes and preference heterogeneity are provided in Section 5.
5 Simulation Results
@@
We simulate [^2] See Appendix G for the specifications of the machine the experiments were run on two scenarios: (1) uniformly random preferences with market sizes $N=K\in\{5,10,15,20\}$; and (2) varying preference heterogeneity at $N=K=10$. Maximum rewards range from $1$ (least preferred) to $K+1$ (most preferred) with constant gaps $\Delta=1$. Player heterogeneity is controlled as in Ashlagi et al. (2017); Liu et al. (2021) via $\beta\in\{0,10,100,1000\}$—larger $\beta$ yields less heterogeneity, with $\beta\to\infty$ implying identical preferences. Specifically, we sample $x_{k}\sim\mathrm{Uniform}(0,1)$ and $\epsilon_{k,i}\sim\mathrm{Logistic}(0,1)$, set $\overline{\mu}_{i}^{k}=\beta x_{k}+\epsilon_{k,i}$, and define $\mu_{i}^{k}=\#\{j:\overline{\mu}_{i}^{j}\leq\overline{\mu}_{i}^{k}\}$ to maintain $\Delta=1$. Hyperparameters are $\lambda=0.9$ and $\kappa=10$ in PCA-SCA. [^3] There are alternative methods for controlling heterogeneity in player preferences, for example notions like uncoordinated/coordinated markets Ackermann et al. (2008). The method we use corresponds to uncoordinated markets when $\beta$ is low. However, coordinated markets are different, with correlation defined on the edges of the matching graph rather than the vertices. We find that the convergence time in these edge-correlated cases is actually comparable to when preferences are uniformly random, compared with node-correlation, where convergence time increases (see Appendix).
@@
5.1 OCA-UCB in the APKP model
Recall that, in this scenario, players must learn both their own and the arms’ preference orderings. We compare OCA-UCB in the APKP model to CA-UCB in the APCK model, running 6000 steps for random preferences and 3000 for varied heterogeneity. Figure 1 (top) shows that OCA-UCB converges to stability—with only modest slowdowns in larger markets—closely matching CA-UCB when preferences are publicly known. The bottom row illustrates that varying player heterogeneity has negligible impact on convergence. Together with our theory, these results imply that private arm preferences pose little extra difficulty for achieving stability when arms themselves know their own orderings.
Figure 1: Left: APCK model with CA-UCB; Right: APKP model with OCA-UCB. Results averaged over 100 runs. Top row: varying market size with uniformly random preferences; bottom row: varying player preference heterogeneity at $N=K=10$. OCA-UCB converges to stability under dual-sided uncertainty, though more slowly due to increased complexity.
5.2 PCA-UCB and PCA-TS in the APU Model
@@
(a) UCB
(b) Thompson Sampling
@@
Figure 2: Results of APU model experiments using (a) UCB and (b) Thompson Sampling to track beliefs. The top row shows uniformly random preferences; the bottom row shows varied player preference heterogeneity ($N=K=10$). Both markets converge to stability in expectation and player regret vanishes; Thompson Sampling achieves faster, smoother convergence.
Using PCA-UCB
Figure 2(a) summarizes the results of PCA-UCB in the APU model. 20,000 time steps were run for random preferences and 10,000 for varied player preferences. The first row shows results for random preferences, while the second row shows results for varied player preferences.
Figures 1 and 2(a) show that the APU model with PCA-UCB stabilizes slower than the APKP model with OCA-UCB. This is because both players and arms need to learn beliefs from scratch. Nevertheless, with random market preferences, both models exhibit similar trends—stabilization probabilities approach 1 and player regret tends to 0—demonstrating that PCA-UCB can achieve decentralized stable matching despite unknown preferences. The second row in Figure 2(a) shows that varying player preference heterogeneity (measured by $\beta$) does not affect convergence, paralleling CA-UCB’s behavior.
Using PCA-TS
@@
Results for when PCA-TS was used in the APU model are shown in Figure 2(b). Like Figure 2(a) (the UCB results), the first row in Figure 2(b) is when the agent preferences are sampled uniformly at random, and the second row is where the players’ preference heterogeneity is varied. Results from PCA-UCB (Figure 2(a)) and PCA-TS (Figure 2(b)) show similarities in market stability and player regret for all market sizes with uniformly random preferences, with larger markets stabilizing slower. We can also see that PCA-TS converges faster and more smoothly than PCA-UCB. We will discuss this particular property in further detail in the next subsection. The results demonstrate that Algorithm 2 can find a stable match in the APU model using both UCB and Thompson Sampling.
Convergence Comparison: UCB vs. Thompson Sampling
@@
When looking at Figures 2(a) and 2(b), we observe that the Thompson Sampling approach for tracking beliefs outperforms UCB in terms of both convergence speed and smoothness. To quantify this behavior, we employ the convergence proxy defined earlier, now comparing the stability graphs of PCA-UCB and PCA-TS. Setting $\mathcal{X}=1000$ and $\theta=90$, we generate the corresponding lines, with the results summarized in Figure 3. Each color corresponds to a different market size, with the solid line representing PCA-UCB and the dotted line representing PCA-TS. Evidently, the lines for Thompson Sampling reach 1.0 (i.e., 100% of the time-steps exhibit stability $90\%$) faster and remain at that level.
@@
Figure 3: Convergence comparison of UCB (solid) and Thompson Sampling (dotted) in the APU model with varying market sizes and uniformly random preferences. Thompson Sampling converges slightly faster and remains stable, while UCB exhibits intermittent dips before ultimately converging.
UCB’s convergence proxy approaches 1.0 more slowly and often dips before recovering, because exploration disrupts players’ belief orderings and leads to unstable matches—especially early on. In contrast, PCA-TS only shows this behavior for $N=20$ and converges more smoothly overall. This suggests that Thompson Sampling is a better method for tracking agent beliefs in the APU model.
6 Conclusion and Future Work
In this paper, we describe algorithms that allow agents to learn their preferences in repeated two-sided matching. Our algorithms are the first to provably converge without explicit communication between agents (the proposers) when arm preferences are not common knowledge (the APKP model) and when arms themselves do not know their preferences (the APU model). We also demonstrate, through simulation, that using Thompson Sampling converges to stability more quickly and smoothly compared to the UCB approach in the APKP model across a range of settings.
Appendix A Full Algorithm for OCA-UCB
Algorithm 1 Optimistic Conflict Avoiding - UCB (OCA-UCB)
1:Input: $\lambda$, $H$ 2:Output: A stable match 3:for $t\in\{1,\cdots,H\}$ do 4:  for $i\in\{1,\cdots,N\}$ do 5:   if $t==1$ then 6:     $A_{i}(t)\leftarrow a_{k},k\sim\text{Uniform}(\left[K\right])$ 7:   else 8:     $D_{i}(t)\sim Ber(\lambda)$ 9:     if $D_{i}(t)==0$ then 10:      $A_{i}(t)\leftarrow$ GetBestArm($p_{i},t$) 11:     else 12:      $A_{i}(t)\leftarrow A_{i}(t-1)$ 13:     end if 14:   end if 15:  end for 16:  for $k\in\{1,\cdots,K\}$ do 17:   $p_{w}\leftarrow$ ResolveRequests($a_{k},t$) 18:   if $p_{w}\neq\emptyset$ then 19:     $r_{p_{w}},r_{a_{k}}\leftarrow$ Sample-Rewards($p_{w},a_{k}$) 20:     Arm-Update-Belief($a_{k},p_{w},r_{a_{k}}$) 21:     Player-Update-Belief($p_{w},a_{k},r_{p_{w}}$) 22:     $L\leftarrow\left\{p_{j}:A_{j}(t)=a_{k}\right\}\setminus\left\{p_{w}\right\}$ 23:     for $p_{l}\in L$ do 24:      UpdateConflictBelief($p_{l},p_{w},a_{k}$) 25:     end for 26:   end if 27:  end for 28:end for 1:procedure GetBestArm ($p_{i}$, $t$) 2:  $S_{i}(t)\leftarrow\emptyset$ 3:  for $k\in\{1,\cdots,K\}$ do 4:   $p_{\bar{p}}\leftarrow$ $p_{j}\text{ s.t. }\bar{A}_{j}(t-1)=a_{k}$ 5:   if $p_{\bar{p}}\in O_{(i)}^{k_{l}}\vee p_{\bar{p}}=\emptyset$ then 6:     $S_{i}(t)\leftarrow S_{i}(t)\cup\{a_{k}\}$ 7:   end if 8:  end for 9:  $k\leftarrow\underset{a_{j}\in S_{i}(t)}{\mathop{\mathrm{arg\,max}}\limits}\left[\text{UCB}_{i}^{j}\right]$ 10:  return $a_{k}$ 11:end procedure
[…]
Appendix B Full Proof of Convergence of OCA-UCB
As Algorithm 1 progresses, each of the players will have complete sets that represent the arms’ true preferences with respect to the other players. In fact, we show theoretically that this approach to the solution does not affect the convergence guarantee of the CA-UCB algorithm in the APCK Model in Theorem 3.
[…]
Recall some terminology. A triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ is inconsistent if $p_{i}$ believes $p_{i}\succ_{a_{k}}p_{j}$ but the opposite is true. Then, let us define a plausible set constructed by $p_{i}$ to be inconsistent with respect to $p_{j}$ and $a_{k}$ if $a_{k}\in S_{i}(t)$, denoted by $\hat{S}_{ij}^{k}(t)$. Finally, let us define an inconsistent plausible set as being inconsequential at time $t$ if $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{i}^{m}:a_{m}\in S_{i}(t)\}\neq k$ i.e. $p_{i}$ does not attempt $a_{k}$ OR if $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{j}^{m}:a_{m}\in S_{j}(t)\}\neq k$ i.e. $p_{j}$ does not attempt $a_{k}$. We begin the proof of the convergence guarantee by first presenting two lemmas:
Lemma B.1.
Proof.
Lemma B.2.
Proof.
First, begin by noting that there are only finitely many inconsistent triplets in any given APKP model. Then, a simple recursive argument will suffice to show that each one gets resolved or becomes inconsequential. The first and second cases from Lemma B.1 are base cases. The first case corrects $p_{i}$ belief altogether, and in the second case the incorrect belief never gets used in any significant way. In the first sub-case of case three, if $a_{k}$ is in-fact $p_{i}$’s best possible achievable arm, then it will continue getting matched with $a_{k}$ regardless of it’s wrong belief (inconsequential w.r.t $a_{k}$). Nonetheless, if $\exists p_{j}$ at $(t+1)$ such that triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ exists, then we can apply Lemma B.1 recursively. And finally, in the second sub-case of case three from Lemma B.1, $p_{i}$ loses the conflict to some other player which implies that some belief gets updated resulting in the resolution of some triplet. However, if $\exists p_{j}$ at $(t+1)$ such that triplet $Q_{i,j}^{k}=(p_{i},p_{j},a_{k})$ exists, then again we can apply Lemma B.1 recursively. ∎
The proof of Theorem 3 is the immediate consequence of lemmas B.1, and B.2. Once all the inconsistent plausible sets are either resolved, or are inconsequential, the best arm picked from each player’s plausible set will be the same as the one picked from CA-UCB’s plausible set. This implies that this approach to updaing beliefs has no effect on CA-UCB algorithm’s guarantee on convergence. ∎
Appendix C Full Algorithm for PCA-SCA
[…]
Appendix D Full Proof of Convergence of PCA-UCB